The larger of 2 numbers is 8 times the smaller. The difference between them is 42. What are the numbers?
Answer by deborabr(181)

the smaller = x
the larger = 8x
8x-x=42
7x=42
x=42/7
x=6
the smaller = 6
the larger = 8*6 = 48